Can you please help me with this problem:
Your friend is planting a circular garden full of different color pansies. Your friend plans to plant five different colors of equal amounts. The garden has a radius of 15 feet. How many square feet of space will each color cover?

Answer by rothauserc(4718)

Area of the circular garden is pi * r^2, where pi = 22/7 and r = 15 feet
:
Area = (22/7) * 15^2 = 707.14284 approx 707 square feet
:
each color covers 707 / 5 = 141.4 square feet
: